# Athena: Enhancing Multimodal Reasoning with Data-efficient Process Reward Models

### v1

@@ 3 Experiments
Table 1: Results on seven multimodal reasoning benchmarks under Best-of-N (N=8) evaluation. $\dagger$ denotes that the results are from [57]. We mark our results and highlight the best result.
@@
Qwen2.5-VL-7B [3] 36.2 68.1 25.4 41.1 21.8 58.0 47.9
Self-consistency [58] 44.7 71.6 28.6 43.7 22.9 60.1 49.5
[…]
@@
InternVL2.5-8B [8] 23.5 64.5 17.0 22.8 9.4 56.2 36.0
Self-consistency [58] 28.4 66.1 21.1 24.7 13.8 57.8 40.2
[…]
@@
Qwen2.5-VL-72B [3] 49.1 74.2 39.3 47.3 35.9 70.2 55.7
Self-consistency [58] 54.8 77.0 43.1 50.8 37.6 71.1 59.6
@@
Table 2: Results on text-only math benchmarks under Best-of-N (N=8) evaluation. $\dagger$ denotes that the results are from [57]. We mark our results and highlight the best result.
@@
Qwen2.5-7B [61] 91.6 75.5
Self-consistency [58] 93.4 79.9
[…]
@@
Qwen2.5-72B [61] 95.8 83.1
Self-consistency [58] 96.0 86.0
[…]
@@
To comprehensively evaluate the proposed method, we assess Athena-ORM and Athena-PRM under the Best-of-N setting on seven multimodal benchmarks and two text-only benchmarks in Sec. 3.1. Furthermore, we evaluate Athena-PRM on VisualProcessBench, comparing its performance with MLLMs as judges and a recent open-source multimodal PRM, VisualPRM-8B [57], as described in Sec. 3.2. We assess the capability of Athena-7B on multimodal math and reasoning tasks in Sec. 3.3. To validate the effectiveness of our designs for Athena-PRM, an in-depth analysis is provided in Sec. 3.4.
3.1 Best-of-N Evaluation
Benchmarks. We evaluate Athena-ORM and Athena-PRM across multiple multimodal mathematical and logical reasoning benchmarks, including MathVista [36], MathVision [54], MathVerse [68], WeMath [44], DynaMath [70], LogicVista [60], and MMMU [65]. To validate the effectiveness of Athena-ORM and Athena-PRM in the text-only scenario, we conduct experiments on the GSM8K [9] and MATH [18] under the BoN settings. Further details can be found in Appendix A.2.
Settings. We employ Athena-ORM and Athena-PRM as the reward model for BoN evaluation, sampling $N=8$ solutions for every problem by default. For decoding, we set the temperature to 0.8 and use nucleus sampling [20] with top-$p$ set to 0.9, and report the average of five runtimes. For multimodal benchmarks, we choose Qwen2.5-VL-7B/72B [3] and InternVL2.5-8B [8] as policy models. [^4] All models used in this paper are “instruct” version. For simplicity, we ignore “instruct” in the text. For text-only datasets, we choose Qwen2.5-7/72B [61] and Ministral-8B [^5] https://huggingface.co/mistralai/Ministral-8B-Instruct-2410 as policy models. Zero-shot and self-consistency [58] serve as our baselines, alongside a recent public multimodal process reward model, VisualPRM-8B [57].
Training details. To train ORMs and PRMs, we fine-tune Qwen2.5-VL-7B on our dataset as introduced in Sec. 2.3. We use the AdamW optimizer [32] with the weight decay as 0 and set the learning rate as $1e^{-6}$ with cosine decay. We train all models with one epoch and set the global batch size to 64. We use DeepSpeed with zero2 [45] and flash-attention-2 [11] to improve training efficiency. To train ORMs, a new special token <step> is added at the end of solutions for ORMs training. This makes a consistent formulation for ORMs and PRMs. For PRMs, <step> is added to separate all reasoning steps and assign labels to corresponding tokens. We use “$+$” and “--” as labels to denote the correctness of each step. We use 8×8$\times$8 × AMD-MI250 GPUs for training and data synthesis.
Results. Table 1 demonstrates that Athena-ORM and Athena-PRM generally enhance the reasoning performance of MLLMs across policy models of different sizes and benchmarks. Notably, Athena-PRM achieves a +10.2 points improvement on the WeMath dataset with Qwen2.5-VL-7B [3] as the policy model, compared to the zero-shot baseline. ORMs also exhibit general improvement across all benchmarks and policy models, although some gains are limited, such as +0.4 points on MathVerse with Qwen2.5-VL-72B [3]. In contrast, PRMs consistently show substantial performance gains over ORMs, illustrating the benefits of fine-grained rewards in test-time scaling.
Table 2 highlights that Athena-PRM enhances reasoning abilities for text-only inputs in both the Qwen2.5 series and Ministral-8B models. Specifically, Athena-PRM achieves +9.9 points improvement for Ministral-8B on the challenging MATH [18] benchmark, demonstrating its effectiveness in text-only scenarios.
@@ 3.1 Best-of-N Evaluation
GPT-4o-mini [41] 53.6 58.9 57.1 56.7 58.5 57.9
GPT-4o [41] 56.3 60.2 59.7 59.0 63.3 60.3
Gemini-2.0-Flash [15] 58.5 60.1 62.8 66.7 58.7 62.3
@@
MiniCPM-V2.6-8B [62] 44.9 50.9 58.9 46.7 57.4 50.4
LLaVA-OV-7B [26] 45.7 43.0 42.2 44.7 52.5 44.4
LLaVA-OV-72B [26] 46.1 48.4 53.0 57.0 57.3 52.3
Qwen2.5-VL-7B [3] 53.1 51.8 47.8 51.3 54.2 51.0
Qwen2.5-VL-72B [3] 59.2 59.0 59.7 62.9 62.3 60.5
InternVL2.5-8B [8] 47.1 45.5 47.8 50.3 50.8 48.0
InternVL2.5-26B [8] 48.8 47.4 49.2 50.4 51.4 49.2
InternVL2.5-38B [8] 51.5 48.4 50.9 51.8 52.5 50.8
InternVL2.5-78B [8] 52.0 51.7 53.7 50.8 52.5 52.6
@@ 3.2 Evaluation on Direct Judgment of Steps
Setup. In addition to Best-of-N evaluation, we assess Athena-PRM on VisualProcessBench [57], comparing it with general MLLMs and VisualPRM-8B [57]. VisualProcessBench [57] aims to evaluate the ability to directly judge the correctness of each step. For MLLMs, we prompt the model to analyze and judge the correctness of each step. We report the F1-score for every subset and the micro-F1 score overall.
[…]
@@ 3.3 Evaluation on the Fine-tuned Model
Table 4: Comparison of multimodal reasoning and mathematical performance across different models. $\Delta$ denotes improvements compared to Qwen2.5-VL-7B [61]. We highlight improvements at least +0.5 points with green.
@@
GPT-4o [41] 72.9 50.6 71.6 43.8 49.9 48.5 64.4
Claude-3.7-Sonnet [1] 71.0 49.3 66.8 41.9 46.7 39.7 58.2
Gemini-2.0-Pro [15] 72.6 56.5 71.3 48.1 67.3 43.3 53.2
@@
InternVL2.5-78B-MPO [56] 68.2 37.6 76.6 36.2 43.7 21.2 50.8
InternVL2.5-78B [8] 70.1 39.8 70.6 32.2 39.2 19.2 49.0
Qwen2.5-VL-72B [3] 68.2 49.1 74.2 39.3 47.3 35.9 55.7
@@
InternVL2.5-8B [8] 56.2 20.2 58.3 20.0 20.4 9.2 33.6
MiniCPM-o-2.6-8B [62] 50.9 25.2 73.3 21.7 35.0 10.4 36.0
Qwen2.5-VL-7B [3] 58.0 36.2 68.1 25.4 41.1 21.8 47.9
@@
Setup. We select Qwen2.5-VL-7B [3] as the initial policy model, obtaining training data as outlined in Sec. 2.4. We use AdamW [32] optimizer and set the learning rate as $1e^{-6}$ with cosine decay. The batch size is set to 128, and the number of epochs is set to one. We denote our fine-tuned model as Athena-7B and evaluate it on the same benchmarks as Sec. 3.1 in the zero-shot setting. For evaluation, we use VLMEvalKit [14] to evaluate Athena-7B and report the results of all baselines from the open leaderboard. [^6] https://huggingface.co/spaces/opencompass/Open_LMM_Reasoning_Leaderboard
Results.   As shown in Table 4, Athena-7B consistently improves upon Qwen2.5-VL-7B [3] across all benchmarks, with significant enhancements of at least $+$0.5 points on five benchmarks. Athena-7B outperforms large open-source models (e.g. InternVL2.5-78B [8]) on five benchmarks. On the MathVista  [36], Athena-7B even achieves the comparable performance with proprietary models such as Gemini-2.0-Pro [15], Claude-3.7-Sonnet [1], and GPT-4o [41].
@@ 3.4 Analysis
Test time scaling. To evaluate the scalability of Athena-ORM and Athena-PRM, we compare their performance using different numbers of samples ranging from 4 to 64 on the MathVista [36]. As shown in Figure 2, reward ranking for solutions at test time improves reasoning performance across different policy models and number of samples. Specifically, ORMs and PRMs improve the self-consistency [58] baseline by 0.8 and 2.1 points with Qwen2.5-VL-72B [3] using eight samples, respectively. As the sample size increases, the performance improves steadily. However, the performance gain of ORMs is markedly lower than that of PRMs. For instance, Athena-PRM achieves 82.4% with Qwen2.5-VL-72B [3] using 64 samples, while Athena-ORM achieves similar performance with self-consistency [58] baseline. These findings underscore the advancements and scalability of PRMs in test-time scaling scenarios.
@@
Figure 2: Best-of-N results on the MathVista [36] across different policies. The number of solutions we sample is from 4 to 64.
Table 5: Ablation study of our design choices on the MathVista [54] and WeMath [44] with Qwen2.5-VL-7B [3] as the policy model.
@@
The effectiveness of our design choices.  Table 5 presents results under various design choices. Our findings indicate that PRMs trained with selectively chosen 5K samples significantly outperform those trained with randomly selected samples, achieving scores 74.1 vs.73.1 and 74.8 vs.73.6 when initialized from ORMs. Moreover, a large-scale dataset of 300K samples labeled using vanilla MC-based methods yields similar performance to our selected 5K samples. With these selected samples, ORM initialization enhances performance by 0.7 points (74.8 vs.74.1), and up-sampling negative labels further boosts performance by 0.4 points on the MathVista [36].
Computation cost analysis.  Vanilla MC-based estimation [55] incurs substantial computational demands to estimate the correctness of intermediate steps. Our prediction consistency filter strategy significantly reduces this burden, requiring only 5K samples for training, thereby lowering the computational cost by a factor of 60. For data synthesis, we employ vLLM primarily [25] to accelerate MLLM inference. Compared to the MC-300K dataset, our Athena-5K dataset reduces GPU hours by approximately 45-fold while achieving similar performance.
4 Related Work
Test time scaling.  While scaling data and model parameters during training has been extensively studied [23, 19], scaling computation at test time has garnered interest more recently [50, 46]. TTS enables language models to leverage additional computational resources when faced with challenging questions. Our study primarily focuses on parallel scaling [67], wherein language models generate multiple outputs simultaneously and aggregate them into a final answer. This aggregation can involve selecting the most common outcome [58] or using reward models [50] to rank solutions, selecting the one with the highest reward, as demonstrated in our Best-of-N settings.
Process reward models.  Reward models are crucial in reinforcement learning [17, 47, 55, 42, 4, 13] and test time scaling [46, 28, 55, 9, 50]. Outcome reward models [9] assign a scalar value to question-response pairs, whereas process reward models evaluate each step, typically achieving superior performance and generalization [28, 37, 46, 27]. While better performance and generalization are achieved, collecting data to train PRMs is challenging. PRM800K [28] is the first open-source process supervision dataset annotated by humans, but scaling it is challenging due to the time and skill required for annotating reasoning steps. Math-Shepherd [55] proposes an automated method for estimating intermediate step correctness using Monte Carlo estimation, though it generates some incorrect labels and demands extensive computational resources. Our work explores strategies to reduce computational costs while improving process label accuracy.

### v2

@@ 3 Experiments
To comprehensively evaluate the proposed method, we assess Athena-ORM and Athena-PRM under the Best-of-N setting on seven multimodal benchmarks and two text-only benchmarks in Sec. 3.1. Furthermore, we evaluate Athena-PRM on VisualProcessBench, comparing its performance with MLLMs as judges and a recent open-source multimodal PRM, VisualPRM-8B (Wang et al., 2025), as described in Sec. 3.2. We assess the capability of Athena-7B on multimodal math and reasoning tasks in Sec. 3.3. To validate the effectiveness of our designs for Athena-PRM, an in-depth analysis is provided in Sec. 3.4.
3.1 Best-of-N Evaluation
Benchmarks. We evaluate Athena-ORM and Athena-PRM across multiple multimodal mathematical and logical reasoning benchmarks, including MathVista (Lu et al., 2024b), MathVision (Wang et al., 2024a), MathVerse (Zhang et al., 2024), WeMath (Qiao et al., 2025), DynaMath (Zou et al., 2025), LogicVista (Xiao et al., 2024), and MMMU (Yue et al., 2024). To validate the effectiveness of Athena-ORM and Athena-PRM in the text-only scenario, we conduct experiments on the GSM8K (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) under the BoN settings. Further details can be found in Appendix C.2.
Settings. We employ Athena-ORM and Athena-PRM as the reward model for BoN evaluation, sampling $N=8$ solutions for every problem by default. For decoding, we set the temperature to 0.8 and use nucleus sampling (Holtzman et al., 2020) with top-$p$ set to 0.9, and report the average of five runtimes. For multimodal benchmarks, we choose Qwen2.5-VL-7B/72B (Bai et al., 2025) and InternVL2.5-8B (Chen et al., 2024) as policy models. [^2] All models used in this paper are “instruct” version. For simplicity, we ignore “instruct” in the text. For text-only datasets, we choose Qwen2.5-7/72B (Yang et al., 2024) and Ministral-8B [^3] https://huggingface.co/mistralai/Ministral-8B-Instruct-2410 as policy models. Zero-shot and self-consistency (Wang et al., 2023) serve as our baselines, alongside a recent public multimodal process reward model, VisualPRM-8B (Wang et al., 2025).
Table 1: Results on seven multimodal reasoning benchmarks under Best-of-N (N=8) evaluation. $\dagger$ denotes that the results are from Wang et al. (2025). We mark our results and highlight the best result.
@@ 3.1 Best-of-N Evaluation
Qwen2.5-VL-7B (Bai et al., 2025) 36.2 68.1 25.4 41.1 21.8 58.0 47.9
Self-consistency (Wang et al., 2023) 44.7 71.6 28.6 43.7 22.9 60.1 49.5
$\text{VisualPRM-8B}^{\dagger}$ (Wang et al., 2025) 39.8 70.3 31.3 44.3 23.0 58.6 48.3
@@
InternVL2.5-8B (Chen et al., 2024) 23.5 64.5 17.0 22.8 9.4 56.2 36.0
Self-consistency (Wang et al., 2023) 28.4 66.1 21.1 24.7 13.8 57.8 40.2
$\text{VisualPRM-8B}^{\dagger}$ (Wang et al., 2025) 36.5 68.5 25.7 35.8 18.0 60.2 43.8
@@
Qwen2.5-VL-72B (Bai et al., 2025) 49.1 74.2 39.3 47.3 35.9 70.2 55.7
Self-consistency (Wang et al., 2023) 54.8 77.0 43.1 50.8 37.6 71.1 59.6
@@
Table 2: Results on text-only math benchmarks under Best-of-N (N=8) evaluation. $\dagger$ denotes that the results are from Wang et al. (2025). We mark our results and highlight the best result.
@@
Qwen2.5-7B (Yang et al., 2024) 91.6 75.5
Self-consistency (Wang et al., 2023) 93.4 79.9
$\text{VisualPRM-8B}^{\dagger}$ (Wang et al., 2025) 94.5 81.6
@@
Qwen2.5-72B (Yang et al., 2024) 95.8 83.1
Self-consistency (Wang et al., 2023) 96.0 86.0
$\text{VisualPRM-8B}^{\dagger}$ (Wang et al., 2025) 96.5 85.2
@@
Training details. To train ORMs and PRMs, we fine-tune Qwen2.5-VL-7B on our dataset as introduced in Sec. 2.4. We use the AdamW optimizer (Loshchilov & Hutter, 2019) with the weight decay as 0 and set the learning rate as $1e^{-6}$ with cosine decay. We train all models with one epoch and set the global batch size to 64. We use DeepSpeed with zero2 (Rajbhandari et al., 2020) and flash-attention-2 (Dao, 2024) to improve training efficiency. To train ORMs, a new special token <step> is added at the end of solutions for ORMs training. This makes a consistent formulation for ORMs and PRMs. For PRMs, <step> is added to separate all reasoning steps and assign labels to corresponding tokens. We use “++” and “-” as labels to denote the correctness of each step. We use $8\times$ AMD-MI250 GPUs for training and data synthesis.
Results. Table 1 demonstrates that Athena-ORM and Athena-PRM generally enhance the reasoning performance of MLLMs across policy models of different sizes and benchmarks. Notably, Athena-PRM achieves a +10.2 points improvement on the WeMath dataset with Qwen2.5-VL-7B (Bai et al., 2025) as the policy model, compared to the zero-shot baseline. ORMs also exhibit general improvement across all benchmarks and policy models, although some gains are limited, such as +0.4 points on MathVerse with Qwen2.5-VL-72B (Bai et al., 2025). In contrast, PRMs consistently show substantial performance gains over ORMs, illustrating the benefits of fine-grained rewards in test-time scaling.
Table 2 highlights that Athena-PRM enhances reasoning abilities for text-only inputs in both the Qwen2.5 series and Ministral-8B models. Specifically, Athena-PRM achieves +9.9 points improvement for Ministral-8B on the challenging MATH (Hendrycks et al., 2021) benchmark, demonstrating its effectiveness in text-only scenarios.
@@
GPT-4o-mini (OpenAI, 2024) 53.6 58.9 57.1 56.7 58.5 57.9
GPT-4o (OpenAI, 2024) 56.3 60.2 59.7 59.0 63.3 60.3
Gemini-2.0-Flash (Gemini Team, 2023) 58.5 60.1 62.8 66.7 58.7 62.3
@@
MiniCPM-V2.6-8B (Yao et al., 2025) 44.9 50.9 58.9 46.7 57.4 50.4
LLaVA-OV-7B (Li et al., 2025) 45.7 43.0 42.2 44.7 52.5 44.4
LLaVA-OV-72B (Li et al., 2025) 46.1 48.4 53.0 57.0 57.3 52.3
Qwen2.5-VL-7B (Bai et al., 2025) 53.1 51.8 47.8 51.3 54.2 51.0
Qwen2.5-VL-72B (Bai et al., 2025) 59.2 59.0 59.7 62.9 62.3 60.5
InternVL2.5-8B (Chen et al., 2024) 47.1 45.5 47.8 50.3 50.8 48.0
InternVL2.5-26B (Chen et al., 2024) 48.8 47.4 49.2 50.4 51.4 49.2
InternVL2.5-38B (Chen et al., 2024) 51.5 48.4 50.9 51.8 52.5 50.8
InternVL2.5-78B (Chen et al., 2024) 52.0 51.7 53.7 50.8 52.5 52.6
@@ 3.2 Evaluation on Direct Judgment of Steps
Setup. In addition to Best-of-N evaluation, we assess Athena-PRM on VisualProcessBench (Wang et al., 2025), comparing it with general MLLMs and VisualPRM-8B (Wang et al., 2025). VisualProcessBench (Wang et al., 2025) aims to evaluate the ability to directly judge the correctness of each step. For MLLMs, we prompt the model to analyze and judge the correctness of each step. We report the F1-score for every subset and the micro-F1 score overall.
Results. The results of VisualProcessBench are presented in Table 3. It is evident that current general open-source MLLMs struggle to judge the correctness of reasoning steps, with most $\scriptstyle\sim$7B models performing no better than the random guessing baseline. Conversely, multimodal process reward models achieve superior performance, rivaling proprietary models. Notably, our process reward model Athena-PRM, with only 7B parameters, achieves a +3.9 points improvement compared to VisualPRM-8B (Wang et al., 2025), even outperforming proprietary models as judges.
@@ 3.3 Evaluation on the Fine-tuned Model
Table 4: Comparison of multimodal reasoning and mathematical performance across different models. $\Delta$ denotes improvements compared to Qwen2.5-VL-7B (Yang et al., 2024). We highlight improvements at least +0.5 points with green.
@@
GPT-4o (OpenAI, 2024) 72.9 50.6 71.6 43.8 49.9 48.5 64.4
Claude-3.7-Sonnet (Anthropic, 2025) 71.0 49.3 66.8 41.9 46.7 39.7 58.2
Gemini-2.0-Pro (Gemini Team, 2023) 72.6 56.5 71.3 48.1 67.3 43.3 53.2
@@
InternVL2.5-78B-MPO (Wang et al., 2024c) 68.2 37.6 76.6 36.2 43.7 21.2 50.8
InternVL2.5-78B (Chen et al., 2024) 70.1 39.8 70.6 32.2 39.2 19.2 49.0
Qwen2.5-VL-72B (Bai et al., 2025) 68.2 49.1 74.2 39.3 47.3 35.9 55.7
@@
InternVL2.5-8B (Chen et al., 2024) 56.2 20.2 58.3 20.0 20.4 9.2 33.6
MiniCPM-o-2.6-8B (Yao et al., 2025) 50.9 25.2 73.3 21.7 35.0 10.4 36.0
Qwen2.5-VL-7B (Bai et al., 2025) 58.0 36.2 68.1 25.4 41.1 21.8 47.9
@@
Setup. We select Qwen2.5-VL-7B (Bai et al., 2025) as the initial policy model, obtaining training data as outlined in Sec. 2.5. We use AdamW (Loshchilov & Hutter, 2019) optimizer and set the learning rate as $1e^{-6}$ with cosine decay. The batch size is set to 128, and the number of epochs is set to one. We denote our fine-tuned model as Athena-7B and evaluate it on the same benchmarks as Sec. 3.1 in the zero-shot setting. For evaluation, we use VLMEvalKit (Duan et al., 2024) to evaluate Athena-7B and report the results of all baselines from Open LMM Reasoning Leaderboard.
Results.   As shown in Table 4, Athena-7B consistently improves upon Qwen2.5-VL-7B (Bai et al., 2025) across all benchmarks, with significant enhancements of at least ++0.5 points on five benchmarks. Athena-7B outperforms large open-source models (e.g. InternVL2.5-78B (Chen et al., 2024)) on five benchmarks. On the MathVista  (Lu et al., 2024b), Athena-7B even achieves the comparable performance with proprietary models such as Gemini-2.0-Pro (Gemini Team, 2023), Claude-3.7-Sonnet (Anthropic, 2025), and GPT-4o (OpenAI, 2024).
@@ 3.4 Analysis
test-time scaling. To evaluate the scalability of Athena-ORM and Athena-PRM, we compare their performance using different numbers of samples ranging from 4 to 64 on the MathVista (Lu et al., 2024b). As shown in Figure 2, reward ranking for solutions at test-time improves reasoning performance across different policy models and number of samples. Specifically, ORMs and PRMs improve the self-consistency (Wang et al., 2023) baseline by 0.8 and 2.1 points with Qwen2.5-VL-72B (Bai et al., 2025) using eight samples, respectively. As the sample size increases, the performance improves steadily. However, the performance gain of ORMs is markedly lower than that of PRMs. For instance, Athena-PRM achieves 82.4% with Qwen2.5-VL-72B (Bai et al., 2025) using 64 samples, while Athena-ORM achieves similar performance with self-consistency (Wang et al., 2023) baseline. These findings underscore the advancements and scalability of PRMs in test-time scaling scenarios.
@@
Figure 2: Best-of-N results on the MathVista (Lu et al., 2024b) across different policies. The number of solutions we sample is from 4 to 64.
Table 5: Ablation study of our design choices on the MathVista (Wang et al., 2024a) and WeMath (Qiao et al., 2025) with Qwen2.5-VL-7B (Bai et al., 2025) as the policy model.
@@
The effectiveness of our design choices.  Table 5 presents results under various design choices. Our findings indicate that PRMs trained with selectively chosen 5K samples significantly outperform those trained with randomly selected samples, achieving scores 74.1 vs.73.1 and 74.8 vs.73.6 when initialized from ORMs. Moreover, a large-scale dataset of 300K samples labeled using vanilla MC-based methods yields similar performance to our selected 5K samples. With these selected samples, ORM initialization enhances performance by 0.7 points (74.8 vs.74.1), and up-sampling negative labels further boosts performance by 0.4 points on the MathVista (Lu et al., 2024b).
Computation cost analysis.  Vanilla MC-based estimation (Wang et al., 2024b) incurs substantial computational demands to estimate the correctness of intermediate steps. Our prediction consistency filter strategy significantly reduces this burden, requiring only 5K samples for training, thereby lowering the computational cost by a factor of 60. For data synthesis, we employ vLLM primarily (Kwon et al., 2023) to accelerate MLLM inference. Compared to the MC-300K dataset, our Athena-5K dataset reduces GPU hours by approximately 45-fold while achieving similar performance.